# An Ising Model of Bursts and Fizzles in Intergroup Violence

### v8

@@ Abstract
During intergroup confrontations, agitating stimuli such as opponents’ threats and provocations can trigger collective violence. We examine video recordings of street fights between groups of young men. Collective violence sometimes breaks out in a burst and at other times in a fizzle wherein only few group members participate. These two patterns are accurately predicted by an Ising model with asymmetric spins.
@@ 1 Introduction
The centennial Ising spin glass model has been applied to a wide range of phenomena [29, 26], among others in social life [1, 19, 7, 10, 32, 15]. One important social phenomenon—cooperation for public goods—has been rarely studied by this model, however, because the symmetry of the spin values (-1 and 1) mismatches the asymmetry of defection ($D$) and cooperation ($C$). Although a complicated workaround is possible for a two-person public goods game [1], we obtain a simpler and more general model by abandoning the symmetry.
Our purpose is to predict and explain when, how, and in what kind of pattern collective action commences, in particular in violent conflicts. In our empirical data, the public good of a group is the victory over, or the defense against, opponents. Our subjects are unarmed and not professionally trained fighters, with limited ability to estimate their payoffs. Collective action in uncertain situations, such as (potentially) violent confrontations, has been explained by models of thresholds [17], cascades [34], and critical mass [23]. These models are built on the assumption that initiative takers or leaders initiate cooperation. Threshold and critical mass models also depend on the assumption that participants have (nearly) perfect information about the numbers of defectors and cooperators. The Ising model does not rely on these assumptions, and is therefore more parsimonious. Individuals only have information about their network neighbors. At a confrontation, focal group members are agitated by opponents’ provocations, threats, and movements—turmoil, in short—that corresponds to temperature in the original model. Due to turmoil, one or few individuals may accidentally start cooperating (fighting, in our case), called “trembling hands” in game theory [12]. This effect happens to be embedded in the Metropolis algorithm [6] used to solve the Ising model [3]. When few individuals start cooperating, they influence their network neighbors, and when turmoil passes a critical level, there is a burst of cooperation. In answering the how question, the Ising model thus points out that cooperation can start by accidental cooperators who are agitated, rather than exceptionally zealous and/or well-informed individuals. If there are initiative takers or leaders, however, they can be accommodated [5].
The novelty of this paper is in its prediction of the temporal unfolding of collective action. When a group consists of unconditional defectors, who are unresponsive to turmoil and cooperating network neighbors, and conditional cooperators [8], who cooperate if they are agitated and/or enough others cooperate, in proportion $p:1-p$, respectively, there is a critical level $p_{c}$ that separates two different patterns. If $p\leq p_{c}$, collective action breaks out in a burst, whereas if $p>p_{c}$, there is no burst but a fizzle of few group members who contribute asynchronously. Furthermore, small groups and small clusters in large groups are more easily agitated to cooperate than large groups, starting at lower levels of turmoil. The predicted proportion, $p_{c}$, and the size effect are strongly supported by video recorded street fights.
2 Ising model
Network ties (coupling) and spins (behavior) are defined differently than in the usual Ising model [29, 26]. Network tie $A_{ij}>0$ means that individual $i$ is in close proximity of and pays attention to group member $j$, or else $A_{ij}=0$. We assume that attention is reciprocal but not necessarily equal. Because attention is scarce and people tend to respond to proportions of their social environment rather than absolute numbers [34], ties are row-normalized, $w_{ij}=A_{ij}/\sum_{j=1}^{n}A_{ij}$, with $n$ the size of the group. A group member can contribute to the public good, with behavioral variable $S_{i}=C$, or defect, $S_{i}=-D$, and $0<D<C$. With row-normalization and asymmetric spins, our Ising model [6] is the Hamiltonian
$H=-\sum_{i\neq j}w_{ij}S_{i}S_{j}.$ (1)
In our examples and empirical study, we choose $C=1$. For the dilemma of cooperation to be nontrivial, $D$ must be chosen such that $H$ has a global minimum when the number of cooperators, $N_{C}=n$, and a single maximum in between $N_{C}=0$ and $N_{C}=n$ (Fig. 1). We chose $D=1/2$ in an earlier paper [6] to maximally distinguish $D$ from the trivial values $D=1$ and $D=0$.
Fig. 1 plots $H/n$ with $N_{C}/n$ for group sizes $n=2$ and $n=5$. The hill is a graphical representation of the dilemma of collective action.
@@ 2 Ising model
Figure 1: The dilemma of collective action presented as a hill in between full defection (left) and full cooperation (right), with the proportion of cooperators $(N_{C}/n)$ on the horizontal axis; $C=1$ and $D=1/2$. The vertical axis $(H/n)$ could be intuited as a negative likelihood; moving uphill from a state of defection $(N_{C}/n=0)$ is very unlikely. A proportion of unconditional defectors, $p$, means that even if all conditional cooperators contribute, $N_{C}/n\leq 1-p$.
We do not assume that individuals know their payoffs in advance, but they will heuristically—and perhaps wrongly—distinguish between valuable ($C>D$) and unvaluable ($C<D$) public goods. Payoffs are defined to facilitate interpretation, but they play no role in solving the model. The payoff for a cooperator is the same as in evolutionary game theory [30]; when an actor cooperates among $N_{C}$ other cooperators, $P_{C}=\theta(N_{C}+1)/n-1$, with an enhancement factor $\theta\geq 1$. A defector’s payoff is almost the same as in game theory, $P_{D}=\theta N_{C}/n+Q$, except for a factor $Q$, which establishes that $P_{D}$ approximates $P_{C}$ when $D$ approximates $C$ [$Q=(\theta/n-1)(1-R);R=(C-D)/(C+D)$; and $\theta=\theta_{0}+R$ with a base rate $\theta_{0}\geq 1$].
The values of $D$ and $C$ can be related to the familiar symmetric Ising model through a mapping
$\{-D,C\}\rightarrow\{S_{0}-\Delta,S_{0}+\Delta\}$ (2)
with a bias $S_{0}=(C-D)/2$ with respect to 0. Then, the two behavioral options become symmetrical with respect to $S_{0}$ at an offset ±$\pm\Delta$. It has been shown that the asymmetry of $S$ is equivalent to the symmetric model ($S_{0}=0$) with an external field $2S_{0}$ [6]. If $\Delta$ is set to a fixed value (0.75 in our examples), decreasing $S_{0}$ makes cooperation less valuable, and is equivalent to an increasing threshold of cooperating network neighbors to win actors over to cooperate, also in other binary decision models [34, 17]. A case in point is an opponent group that is (much) larger than the focal group. Cooperation becomes more appealing when $S_{0}$ increases, which corresponds to a decreasing threshold or proportion of cooperating network neighbors. This happens, for example, when group members overestimate their fighting skills.
The dynamics of a group are driven by agitating stimuli, $T$, analogous to temperature in the original model. In our empirical study, $T$ expresses the (aggregated) effect of opponents on members of a focal group. Initially, all focal group members defect. If $p=0$ and $T$ increases past a critical level $T_{c}$, there is a burst of cooperation wherein the majority of group members participates (Fig. 2). If $p$ increases, $T_{c}$ increases, and the proportion of participants in the burst ($N_{C}/n$) decreases; if $p>p_{c}$, the bursty pattern does not appear. Fighting ends when exhaustion sets in, a winner stands out, or others intervene. To determine $p_{c}$, we will conduct a mean field analysis.
@@
Figure 2: (Color online.) Proportion of cooperators ($N_{C}/n$) with agitation ($T$) in a fully connected network in the size range of our empirical study ($n=5$); $C=1$ and $D=1/2$. The black line depicts the group without unconditional defectors ($p=0$), the red line with one unconditional defector ($p<p_{c}$), and the dashed blue line with two unconditional defectors ($p>p_{c}$).
3 Mean field analysis
For the mean field analysis, the proportion of cooperators, $N_{C}/n$, is expressed in terms of the order parameter, $M=1/n\sum_{i=1}^{n}S_{i}$. Consequently, $N_{C}/n=(M+D)/(C+D)$.
The mean field assumption can be stated as $S_{i}=\bar{S_{i}}=M$. Starting out with the Hamiltonian (Eq. 1), using the mapping from Eq. 2, taking the row-normalization of the adjacency matrix into account, and leaving out the Boltzmann constant, it has been established [6] that
$M=S_{0}+\Delta\tanh{(\Delta M/T)}.$ (3)
Here, we distinguish unconditional defectors, in proportion $p$, from conditional cooperators, $1-p$. Accordingly,
$M=S_{0}-p\Delta+(1-p)\Delta\tanh{(\Delta M/T)}.$ (4)
Identifying critical $T$ (see Supplementary Material) leads to
$-\frac{\operatorname{arcosh}(\sqrt{\Delta^{2}(1-p)/T})}{\Delta^{2}/T}+$
$(1-p)\sqrt{1-\frac{1}{\Delta^{2}(1-p)/T}}=\frac{S_{0}}{\Delta}-p.$ (5)
Solutions of this equation become complex if $p>\frac{S_{0}}{\Delta}$, hence $p\leq\frac{S_{0}}{\Delta}$. The choice of $C=1$ and $D=1/2$ implies that there is no (burst of) cooperation if $p>1/3$. We are now to find out what the mean field result tells us about actual street fights.
4 Empirical study
Bystanders often film street violence with their phones, and upload their videos to websites such as YouTube, LiveLeak, and WorldStarHipHop, where we collected our data. Our sample from these sites is not random with respect to violence, hence we cannot test if turmoil is its cause, but in all likelihood it is random with respect to the temporal patterns of violence. We analyzed 59 groups in the size range $102\leq n<$ 10 (and one larger group, with six active members); see Supplementary Material for details and examples. The videos were coded using Noldus Observer XT 14 software. Videos were played at half speed many times over. Each of 406 individuals was coded for belonging to a focal, opponent, or third-party group. Their behavior was interpreted and represented on the timeline.
4.1 Operationalization

### v9

# Intergroup violence in bursts or fizzles

@@ Abstract
During intergroup confrontations, agitating stimuli such as opponents’ threats and provocations can trigger collective violence, even when the usual mechanisms of cooperation, such as norms with sanctions, are absent. We examine video recordings of street fights between groups of young men. Collective violence in their attacks sometimes breaks out in a burst, while at other times it reaches only a fizzle in which only a few group members participate. An adapted Ising spin-glass model demonstrates that these two temporal unfoldings can be predicted by the proportion of unconditional defectors in a focal group.
@@ 1 Introduction
When observing the unfolding of intergroup violence committed by non-professionals, one notices a burst or a fizzle in the majority of cases. Before punches are thrown, individuals in the focal group and their opponents insult and threaten, and thereby agitate, one another. For this to result in violence by focal group members, they have to be in close proximity and have to have face-to-face contact [15, 14, 36]. When agitation passes a critical level (for example, when one individual is pushed by an opponent), violence can break out in a burst wherein a majority of individuals (or a subgroup from a larger and more dispersed group) participates. Alternatively, the confrontation can fizzle out, even if group members are agitated. In the case of fizzling, one or few individuals push or punch asynchronously, but there is no burst of violence. Additionally, understanding how collective violence unfolds is complicated by other people present on the scene, who may form an audience or try to de-escalate [32, 46, 63]. De-escalation happens often in street fights [47]. Given a common prelude of proximity and agitation mixed with de-escalation, our question is how to explain bursts and fizzles. To this end we use the centennial Ising spin-glass model from physics [55, 44], which resulted in a Nobel Prize for Parisi (2021), and has been used occasionally in the social sciences in various studies of social influence [12, 54, 29, 22]. We apply the model to videos of street fights between groups of (mostly) young men.
Collective violence involves a dilemma of collective action [26, 43, 42], so before predicting bursts and fizzles, we must first address how the Ising model explains overcoming such dilemmas. The face-to-face contacts of the focal group members are the social ties of their network. We already know from public goods experiments that more than half of the participants are conditional cooperators willing to contribute if others do [13], thus conforming to their (weighted) average neighbor in the network [7, 57]. Our empirical study is about relatively inexperienced individuals, rather than seasoned fighters, hence our subjects experience uncertainty when they come under threat, which increases the likelihood of their conforming to the group [65, 39]. Conformity makes sense from an evolutionary perspective when payoffs are hard to predict [59]. Therefore subjects’ intentionality not only depends on their expected payoff (as a share of the public good, in this case attack of or defense against opponents) but also their desire to align with their group members.
Now think of two individuals, each with the behavioral options to defect or cooperate. (Behavioral options correspond to magnetic spins in the original model, but that does not matter here.) The possibilities are: (1) both individuals defect, which avoids exploitation but does not yield any public good; (2) one individual freerides the cooperating other, yielding half of the public good; or (3) both cooperate at a cost, which maximizes the public good. The dashed line in Fig. 1 plots these possibilities, with the normalized number of cooperators ($N_{C}/n$) from left to right. At the beginning, both defect (the local minimum of $H$ on the left). The vertical axis ($H$, or energy in the original) conveys that moving uphill (i.e., that one of the two will start contributing to the public good) is unlikely. By contrast, if moving to lower levels of $H$ is possible, it will certainly happen; thus, if one of the two cooperates, the other conditional cooperator will join in, and the curve goes to the global minimum of $H$, on the right. The hill is the graphical representation of the dilemma of collective action; it is also drawn for a group of five. It shows that without a driving force, it is very unlikely that a group of non-fighting individuals ($n\geq 2$) will overcome the hill and fight collectively, and it is even less likely in larger groups.
@@
Figure 1: The dilemma of collective action presented as a hill between full defection (left) and full cooperation (right), with the proportion of cooperators $(N_{C}/n)$ on the horizontal axis. The vertical axis $(H/n)$ could be intuited as a negative likelihood; moving uphill from a state of defection $(N_{C}/n=0)$ is very unlikely. One line (dashed) is drawn for a dyad and one for a clique of five individuals. The larger the group is, the more rounded the hill becomes.
Nonetheless, some individuals may start fighting, which in turn may prompt others to join. When agitating stimuli, such as opponents’ threats, increase, one or few individuals may accidentally cooperate. This phenomenon of accidental, or spontaneous, cooperation—called “trembling hands” in game theory [17]—can influence proximate others and may result in a cascade of cooperation. For the model, we define turmoil ($T$, or temperature in the original) as opponents’ behavior that agitates members of the focal group. If turmoil increases beyond a critical level ($T_{c}$), fighting can break out in a burst. [^1] The interesting question of why some people self-select into groups where they can fight against others is beyond the scope of this study. Our contribution is to show that the outbreak of collective violence depends on a second critical level, namely the critical proportion of unconditional defectors in the focal group ($p_{c}$). In violent situations, people may defect for several reasons. They may be too scared to fight [15], have empathy with their opponents, disagree with violence (i.e., appraise the public good differently), try to de-escalate, be stopped in their tracks by de-escalators (or guardians), or they may have fought but were wounded, were wrestled to the ground, or became exhausted at some point and became passive. If, for whatever reason, their proportion stays below the critical level ($p\leq p_{c}$), turmoil ($T\geq T_{c}$) will be followed by a burst of violence, but above the critical level ($p>p_{c}$), there is neither a critical level of turmoil nor a burst, and violence occurs in a fizzle. The critical level of steady defectors does not follow from earlier theories, but the Ising model allows us to infer it, which we use as a prediction that we test empirically. Note that $p$ incorporates not only de-escalators in the focal group, but also focal group members who have been prevented from fighting by de-escalators in the focal group or by others.
Most models in the literature revolve around individual rewards and punishments, called selective incentives [42], on top of a share of the public good. These incentives require norms about (in)appropriate behavior in a given situation [19], as well as monitoring of group members [48], and transmission of information (i.e., gossip) through the group’s network, which leads to reputations [41] that feed back through selective incentives, with or without leaders. This package of mechanisms is crucial for ongoing cooperation in the longer run but needs time to develop, which may not be available when threats are imminent. The more time people have, the better they can prepare themselves, which is especially important for high-risk situations. Examples of well-prepared groups are police, soldiers, firefighters, and combat medics who receive professional training that enables them to cooperate effectively and respond to situational stimuli in predetermined manners rather than spontaneously. By contrast, ordinary citizens who face violent opponents are far less prepared, or not prepared at all. For unprepared groups, the uncertainties (of outcomes, benefits, and costs) are higher.
Some earlier models of cooperation do not feature the usual package of cooperation mechanisms; namely, models of thresholds [25], cascades [61], and critical mass [35]. They draw on people who take the initiative or leaders to set cooperation in motion. They also draw on strong rationality assumptions, such as perfect information on the numbers of cooperators. It has been found that in protests, however, leaders often proscribe violence, and most violence occurs in the absence of leaders [27]. Hence, the presence of leaders rarely explains the outbreak of violence. The Ising model, in contrast, points out how cooperation can start spontaneously without leaders, galvanized by accidental cooperators rather than exceptionally zealous ones. [^2] This was already known for the prisoners’ dilemma [17]. If there are initiative takers or leaders [23], however, they can be accommodated, and people may also enter and leave a focal group. The Ising model is also more parsimonious than the aforementioned models because it does not rely on assumptions of strong rationality; the degree of rationality is bounded in violent and other uncertain situations due to incomplete information about opponents and group members beyond face-to-face contact.
2 Model
Members of a (fledgling) focal group, indexed $i$ or $j$, can defect, $D$, or contribute, $C$, to a public good, with $0<D<C$. Behavioral variable $S_{i}$ of conditional cooperators can take the value $S_{i}=C$ or $S_{i}=-D$, whereas unconditional defectors stay put at $S_{j}=-D$. Before a collective action, everyone in a focal group defects. Network ties among focal group members, $A_{ij}>0$, mean that $i$ is in close proximity to, and senses the behavior of, group member $j$, or else $A_{ij}=0$. Because people tend to respond to proportions of their social environment rather than absolute numbers [61, 21], ties are row-normalized, with $w_{ij}=A_{ij}/\sum_{j=1}^{n}A_{ij}$ such that $\sum_{j=1}^{n}w_{ij}=1$. We assume that sensing is reciprocal to at least some degree but not necessarily symmetrical, and that in our small groups ($n<10$), everyone senses all others, thus the network is fully connected. [^3] Temporary exceptions to full connectedness were groups where some people’s view was blocked by objects, opponents, or bystanders, as well as one larger group ($n=14$). These cases we investigated through simulations. Opponents’ turmoil, represented by $T$, has an effect on the focal group at the aggregate level and affects everyone equally. Locally varying turmoil is discussed elsewhere [9]. The Ising model is the following Hamiltonian equation [5, 60]
$H=-\sum_{i\neq j}^{n}w_{ij}S_{i}S_{j}.$ (1)
We do not assume that individuals know their payoffs in advance, but they will heuristically—and perhaps wrongly—distinguish between valuable ($C>D$) and nonvaluable ($C<D$) public goods or, to the same effect, between potentially efficacious and non-efficacious actions [49]. Note that payoffs are not used in the model’s calculations but are defined to provide a meaningful interpretation. When $i$ chooses between $C$ and $D$ amid $N_{C}$ cooperators, payoffs for cooperation and defection are, respectively
$P_{C}$ $=\theta(N_{C}+1)/n-1$ (2)
$P_{D}$ $=\theta N_{C}/n+Q$ (3)
[…]
For our empirical study, we have to choose values for $C$ and $D$ to predict the critical level of unconditional defectors, $p_{c}$. The most obvious choice is $C=1$, as in game theory and the original Ising model. For $D$ we want to avoid two trivial values: if $D=1$, there is no point in cooperating, and if $D=0$, there is no dilemma (but a downward slope to the right in Fig. 1). Choosing $D$ to be maximally distant from the trivial values seems to be a reasonable first approximation, hence we set $S=\{1,-1/2\}$ for all conditional cooperators, also in the examples. Because these two values were published earlier [10], one could say they were preregistered. For unconditional defectors we set $S_{j}=-1/2$, irrespective of their reasons. Note that all earlier Ising models had either the values $\{1,-1\}$ [12, 54, 22, 64] or $\{1,0\}$ [16]. Of all these models, only one represents a public goods game, in this case for two individuals [1, 50] whereas our model is applicable to groups of any size.
[…]
Solving the Ising model boils down to minimizing $H$, which can be done analytically by a mean field approach (in the Supplementary Material). It abstracts away from the network structure, which is no problem for our fully connected networks, and focuses on the mean behavior. We prove that, if $C=1$ and $D=1/2$, $p_{c}=1/3$. For inhomogeneous (such as clustered) networks, the mean field approach is inaccurate. $H$ can also be minimized computationally, which is much simpler, and can take inhomogeneous networks into account. Here, we explain the computational approach (Fig. 2). For a given level of turmoil, a network node $i$ is randomly picked and $H$ is calculated. For comparison, $i$’s current behavior is flipped, from $D$ to $C$ or the other way around, and $H^{\prime}$ with the flip is calculated. The flip is accepted and implemented if $H^{\prime}<H$ or with a certain chance that increases with $T$. In other words, increasing $T$ increases the amount of randomness (due to situational turmoil) in $i$’s decision. A behavioral change of $i$ affects others in $i$’s neighborhood when it is their turn to decide. Consecutive decisions are Monte Carlo steps in the Metropolis algorithm [5] that loops through great many Monte Carlo steps (counted by $t$ in Fig. 2) in order to allow individuals’ interdependent behavior to settle down. This procedure is repeated at increments of $T$, in our study with step size 0.01.
Figure 2: The Metropolis algorithm. Index $i$ runs over the conditional cooperators (not the unconditional defectors); $t$ is a counter of Monte Carlo steps; $c_{r}$ is a random number in the range $0\leq c_{r}\leq 1$.
@@ 2 Model
Figure 3: Color online. Proportion of cooperators ($N_{C}/n$) with turmoil ($T$) in a fully connected network in the size range of our empirical study ($n=5$); again, $C=1$ and $D=1/2$. The black line (top) depicts the group without unconditional defectors ($p=0$), the red line (middle) a group with one unconditional defector ($p<p_{c}$), and the dashed blue line (bottom) a group with two unconditional defectors ($p>p_{c}$).
To illustrate the emergence of cooperation with increasing turmoil in a group of $n=5$, we apply the Metropolis algorithm Fig. 3. First, there are only conditional cooperators. At low turmoil, collective action does not start (black line at the top), but at a critical level $T_{c}$, (almost) everybody bursts into cooperation, with a maximum (where $N_{C}/n\approx 1$) at or near $T_{c}$. Then, a few accidental cooperators win over most others to join the collective action. The effect of turmoil is nonmonotonic and the level of cooperation decreases if $T$ keeps increasing beyond $T_{c}$, which means that very strong turmoil becomes more confusing than agitating. Fighting ends when exhaustion sets in, a winner stands out (opponents have been wrestled to the ground or have fled), or others intervene.
If there are unconditional defectors, $T_{c}$ increases, which is hardly visible in Fig. 3 but more pronounced in larger networks, and maximum cooperation is lower than in the previous case because the number of conditional cooperators is lower. [^6] In Fig. 1, a proportion of unconditional defectors, $p$, means that no matter how many conditional cooperators contribute, $N_{C}/n\leq 1-p$. If the proportion of unconditional defectors surpasses a critical level, $p_{c}$, there is a fizzle (dashed blue line in Fig. 3) with much lower levels of cooperation, and no burst-like start of cooperation. In the mean field analysis, there is no (burst of) cooperation if $p_{c}>S_{0}/\Delta$, independent of network size and density. In simulations, $p_{c}$ is proportionally less precise in smaller networks (Fig. 4; Table S3), and the prediction for the triad (red dot in Fig. 4) is empirically false whereas the mean field is correct: one unconditional defector does not prevent cooperation of the other two. Also one empirical group of six with two unconditional defectors had a burst, which is consistent with the mean field but did not happen in the simulations. Based on the mean field analysis ($p_{c}=1/3$) and a nearly identical result from the simulations ($p_{c}=0.34$), [^7] If unconditional defectors are clustered together in a sparse network, they are less in the way of collective action of the remainder network (thus $p_{c}$ is higher) than if they are evenly spread out across the network. our main prediction is that $p_{c}=1/3$.
Figure 4: For simulated networks with sizes $3\leq n\leq 100$, the minimum number of unconditional defectors who prevent a burst is plotted. The red dot (left, at the bottom) marks the triad (1 defector prevents a burst of 2), which conflicts with the empirical data (1 defector and a burst of 2).
The critical threshold of turmoil increases with network size at a decreasing rate [9], but it also increases with the proportion of unconditional defectors. At $T_{c}$, simulations point out that cooperation starts in small clusters of conditional cooperators, which we will also test empirically. [^8] This bottom up mounting of cooperation is similar to bottom up synchronization [3]. This outcome is puzzling in real life because larger groups have a better chance to win at lower individual costs. Yet if in a small (sub)group, someone starts fighting, he accounts for a relatively large proportion of his neighbors’ social contacts, and more readily wins them over to fight than in a large group where he would comprise a small minority.
3 Data and Methods
In studying violence, lab experiments lack the turmoil and emotional intensity of real violent confrontations due to their obligation to meet ethical standards. Field studies, in contrast, cannot be based on a random sample of participants or groups, yet they are invaluable for realistic view on violence [4]. We obtained 42 videos from websites such as YouTube, LiveLeak, and WorldStarHipHop using search terms with the English keywords “brawl,” “street fight,” and “assault.” This sample is random with respect to temporal unfolding and (sub)group size. Of these clips, 36 are from English-speaking countries (mainly the US and the UK, with one from Canada and one from India); five of the remaining clips are from the Netherlands, and one is from Colombia. We did not observe differences in relevant behavior related to the location of the recording. To keep distracting factors away from our analysis, we excluded clips with professional fighters, long range weapons, protective clothing, a referee, ambush attacks, or youths in a school yard. Most of our videos are recorded on phones by bystanders and are left-truncated; in all likelihood, there would have already been some turmoil that motivated bystanders to start filming. The shortest lasted 30 sec. and the longest was nearly 5 minutes (mean 101 sec.; s.d. 59 sec.). Out of a potential 2 x 42 groups, where the opponents in one analysis become the focal group in the next, 25 groups attacked a single individual rather than a group. Because a lone individual is unable to act collectively, this leaves 59 groups to examine; see the Supplementary Material for all case studies and elaborate examples of interpreting and coding. One group had 14 members, but all other groups were small, $102\leq n<$ 10 (mean 3.6). The smaller groups were simulated as cliques wherein everyone could see one another unless there were obstacles or de-escalators obstructing visual contact.
The videos were coded using Noldus Observer XT 14 software. Clips were played at half speed many times over, and one of us discussed the coding of each with one or two assistants. The assistants were unaware of the theoretical expectations. Each of the 406 individuals was coded for group belonging, and their behavior was interpreted and coded on the timeline.